OpenStudy (gabylovesyou):

Match each shape with its surface area. (Round if necessary) square pyramid: base length 6 meters, slant height 10 meters cone: radius 4 meters, slant height 8 meters square pyramid: base length 5 meters, slant height 11meters cone: radius 6 meters, slant height 4 meters A. 188m^2 B. 135m^2 C. 156m^2 D. 151m^2

OpenStudy (saifoo.khan):

Do you know formulas for each?

OpenStudy (gabylovesyou):

mmmmm yes wait lol

OpenStudy (gabylovesyou):

ohhh no i dont sorry i have volume but not surface area so idk

OpenStudy (gabylovesyou):

do u know?

OpenStudy (saifoo.khan):

Cone: http://0.tqn.com/d/math/1/0/t/L/conerr.jpg

OpenStudy (saifoo.khan):

Pyramid: http://0.tqn.com/d/math/1/0/H/F/sbpyramidr.gif

OpenStudy (anonymous):

SA = s2 + 2 × s × l : pyramid s=base length l=slant height just substitute the given value

OpenStudy (anonymous):

s^2

OpenStudy (gabylovesyou):

SA = 10^2 + 2 x 10 x 6??? Am i right for number 1??

OpenStudy (anonymous):

nope 10 is the slant height

OpenStudy (anonymous):

SA = s2 + 2 × s × l : pyramid s=base length l=slant height just substitute the given value

OpenStudy (anonymous):

@Gabylovesyou

OpenStudy (gabylovesyou):

ohh

OpenStudy (anonymous):

getss??

OpenStudy (gabylovesyou):

so then its SA = 6^2 + 2 x 6 x 10

OpenStudy (anonymous):

yup xD gaby

OpenStudy (gabylovesyou):

156?

OpenStudy (anonymous):

yup \[SA=pi rs +pi r^2\] for cone r=radius s=slant height substitute the given values @Gabylovesyou

OpenStudy (gabylovesyou):

for the other square pyramid question is it 135??

OpenStudy (anonymous):

yup xD nice

OpenStudy (gabylovesyou):

yaay ok! im going to do the cone now

OpenStudy (gabylovesyou):

SA = 3.14 x 4 x 8 + 3.14 x 4^2

OpenStudy (gabylovesyou):

am i right?

OpenStudy (gabylovesyou):

@traile

OpenStudy (anonymous):

yup but it is better if you will use the pi in the calculator 

OpenStudy (gabylovesyou):

pi = 3.14

OpenStudy (anonymous):

yup 3.1416 xD dont forget to use parenthesis

OpenStudy (gabylovesyou):

I got 150.72. Since I don't have that i have to round so i get 151??

OpenStudy (anonymous):

yup

OpenStudy (gabylovesyou):